How much heat (in joules) is used to heat a 53.30 gram sample of iron from 14.7 degrees Celcius to 100.00 degrees Celcius if the specific heat of Fe is 0.450 J/g°C?  J
Answer #1

mass of fe=m= 53030 gm Temperature charge (AT) = 100-14-7= 85-3°c specific heat of fe = c = 0.450 a heat = mcat = 53:304 0:45